OpenStudy (anonymous):

Shelly delivers the weekly local paper to neighborhoods in her town. House numbers are even on one side of the street and odd on the other. Shelly delivers an equal number of papers to both sides of the street. Although she always aims for the front doorstep, Shelly typically misses on three of the tosses on her route each week. Design and conduct a simulation to estimate the probability that next week, Shelly's three misses will all be at odd-numbered houses.

OpenStudy (anonymous):

Hint: You can set up the experiment using 3 coins to collect the data. Allow one side of the coin to represent Heads (evens) and one other side to represent Tails (odds). a) Explain clearly your design of the simulation, including choice of probability tool and description of a single trial. b) Conduct the simulation with trials and record the results. c) Calculate the experimental probability that all 3 of Shelly's missed papers will be at odd-numbered houses.

OpenStudy (anonymous):

So we say that the probability that she misses an odd is 1/2 and the probability that she misses an even is also 1/2. The probability of her missing at 3 odd number houses is (1/2)^3, because they must all happen so they depend on each other. This means multiply the probabilities.

OpenStudy (lukebluefive):

The probability that she will miss throw on one of the odd-numbered houses is: \[\frac{1}{2}\] If she misses on either side with equal probability, the likelihood that she will miss all three on the same side will be: \[\frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} = \frac{1}{8}\]

OpenStudy (anonymous):

Jaycee follows baseball closely. This season, his favorite professional player has a batting average of 0.250 , or 25% . This means that, on average, he gets one hit every four times he is at bat. Jaycee wants to know the likelihood that his favorite player will get at least two hits in the five times he'll be at bat. So Jaycee conducted a simulation to find out. He programmed a random number generator to give numbers between 1 and 4, with an outcome of 1 indicating a hit. His results are shown here. Based on Jaycee's data, what is the probability that his favorite baseball player will get at least two hits in the next game, with five times at bat?

OpenStudy (anonymous):

@AccessDenied